OpenStudy (darkprince14):

f(x) = x/(x^2 - 2x +2) Obtain the power series representation of f in x-1

OpenStudy (experimentx):

you might start with factorizing the denominator (both has complex roots) and then making separate fractions (partial fractions) and then expand using geometric series for 1/((x-1) + something) ... good luck.

OpenStudy (anonymous):

Write it like \[ \frac{x}{x^2-2 x+2}= \\ \frac{x}{x^2-2 x+1+1}=\\ \frac{x}{(x-1)^2+1}=\\ \frac{x-1+1}{(x-1)^2+1}=\\ \frac{x-1}{(x-1)^2+1}+\frac{1} {(x-1)^2+1} \] Use geometric series to expand the last line

OpenStudy (anonymous):

\[ \Large \frac 1{1+u}=\sum_{n=0}^\infty (-1)^n u^n\\ \Large \frac 1{1+u^2}=\sum_{n=0}^\infty (-1)^n u^{2n}\\ \Large \frac 1{1+(x-1)^2}=\sum_{n=0}^\infty (-1)^n (x-1)^{2n}\\ \Large \frac {x-1}{1+(x-1)^2}=\sum_{n=0}^\infty (-1)^n (x-1)^{2n+1}\\ \] Convergence is for \[ |x-1|<1 \]

OpenStudy (anonymous):

Combining, you obtain \[ \frac{x}{x^2-2 x+2}=1+(x-1)-(x-1)^2-(x-1)^3+(x-1)^4+(x-1)^5-(x-1)^6-\\(x-1)^7+(x-1)^8+( x-1)^9+O\left((x-1)^{10}\right) \]

OpenStudy (anonymous):

@experimentX

OpenStudy (anonymous):

Rich Gang Tapout Go Watch On Youtube